The cheetah can reach a top speed of 114 km/h (71 mi/h). While chasing its prey in a short sprint, a cheetah starts from rest and runs 45 m in a straight line, reaching a final speed of 72 km/h.


(a) Determine the cheetah’s average acceleration during the short sprint, and


(b) find its displacement at t = 3.5 s.


1
Expert's answer
2020-10-12T07:44:21-0400

Given:

"v_i=0\:\rm m/s"

"v_f=72\:\rm km/hr=20\:\ m/s"

"d=45\:\rm m"

"t=3.5\:\rm s"

(a) the cheetah’s average acceleration

"a=\frac{v_f^2-v_i^2}{2d}=\rm\frac{20^2-0^2}{2\times 45}=4.44\: m/s^2"

(b) the cheetah’s displacement at time "t"

"d=\frac{at^2}{2}=\rm \frac{4.44\times 3.5^2}{2}=27.2\: m"